What is the management fee (1) on a per available room basis and (2) as a percentage of total revenue for a 255-room hotel located in California that had an occupancy level of 62%, ADR of $84.53, a room revenue to total revenue % of 56.4%, and a gross operating profit % of 24.8%? The management fee agreement stipulated that the company would receive 3% of gross revenue, and 10% of gross operating profit.

1.Please use the information from Question 10 to calculate the Total mgmt fee.

               Total mgmt fee (round to whole number) $ ___

2.Please use the information from Question 10 to calculate the Mgmt fee on PAR basis.

            Mgmt fee on PAR basis (round to two decimal places) $ ___ PAR/yea
Answer #1

Answer:

By using given data,

Total Available rooms in Hotel = 255

Occupancy level = 62%

Average Daily Revenue=$84.53

A Room Revenue/Total Revenue=56.4%

Operating Profit=24.80%

1) Total Management fee:

Total Revenue

Averag room occupied anytime=0.62 * 255 = 158.10 =158

rooms let's say:

Now ADR per room =$84.53

Now total Revenue per day per room=$84.53/(0.564)=$149.8759

Now Managemnet fee is 3% of Gross Revenue= $149.8759 *.03 = $4.496277 and,

10% of gross operating profit=$149.8759 x 0.248 X 0.10=$3.716922

Total Management fee per room=$4.496277 + $3.716922 =$ 8.213199

Total Management fee for the business with 62% occupancy=158*x$ 8.213199=$1,297.685

Managemnt fee as percentage of Total Revenue =$1,297.685/($149.8759*158)=5.48%

2) Now, total Managemnt fee on PAR basis = $1,297.685 / 255 = $5.088962

If you do not get anything in this solution, please put a comment and I will help you out.
Do not give a downvote instantly. It is a humble request. If you like my answer, please give an upvote.